OpenStudy (anonymous):

a body is thrown up from a tower if the ratio of distance to displacement is n the ratio of time for upward and downward motion is?

OpenStudy (shubhamsrg):

do you have a figure?

OpenStudy (anonymous):

no i will give the answer \[\sqrt{n-1}/\sqrt{n+1}\]

OpenStudy (anonymous):

erm i got \[\sqrt{n-2} /\sqrt{n-1}\] will post my workings soon.

OpenStudy (anonymous):

I figured out my mistake shall try again.

OpenStudy (anonymous):